complex numbersson a) Express Z as a complex number in rectangular form. Z = (5 + 12j).(12 + 5j). e 10 b) Express Z as a complex number in p
Answer #1

(a) Z = (5+12j)(12+5j)ej10\pi = (169j)*(cos(10\pi) + jsin(10\pi)) = 169j = 0 + 169j

(b) Z = \frac{2+2j}{2-2j} + 2\angle 45^{o} = \frac{8j}{8} + 2(cos(45o) + jsin(45o)) = j +\sqrt{2} + j\sqrt{2} = \sqrt{2} + j(1+\sqrt{2}) = 2.7979\angle 59.6388^{o}

(c) \frac{200\angle 96^{o}}{10\angle 36^{o}} = 20\angle 60^{o} = 20(cos(60o) + jsin(60o)) = 10 + j17.3205

thus R = 10

L = 17.3205/100 = 0.173205